OpenStudy (anonymous):

Solve 2/5x = -3/15 + 1/x

OpenStudy (anonymous):

the options are 1) x = 3 2) x = 0 3) x = 5 4) x = 15

OpenStudy (anonymous):

\[ \frac { 2} { 5x} = \frac { -3}{ 15} + \frac { 1} { x} \\ (5x)\frac { 2} { 5x} =(5x) \frac { -3}{ 15} + (5x) \frac { 1} { x} \\ 2 = -x + 5\\ x= 3\]

OpenStudy (anonymous):

2/5x=-3/15+1/x put 1/x=y 2y/5=-3/15+y 6y=-3+15y 15y-6y=3 9y=3 y=3/9=1/3 1/x=1/3 x=3

OpenStudy (anonymous):

ok thanks